Gilder Corporation makes a product with the following standard costs:

Standard Quantity or Hours Standard Price or Rate Standard Cost Per Unit
Direct materials 6.10 grams $ 5.00 per gram $ 30.50
Direct labor 0.20 hours $ 11.00 per hour $ 2.20
Variable overhead 0.20 hours $ 4.00 per hour $ 0.80


The company reported the following results concerning this product in June:

Originally budgeted output 3,100 units
Actual output 3,180 units
Raw materials used in production 18,508 grams
Purchases of raw materials 21,108 grams
Actual direct labor-hours 560 hours
Actual cost of raw materials purchases $99,210
Actual direct labor cost $5,757
Actual variable overhead cost $1,594


The company applies variable overhead on the basis of direct labor-hours. The direct materials purchases variance is computed when the materials are purchased.

The materials quantity variance for June is:
Answer #1

Materials Quantity Variance = (SQ-AQ) * SP

= (18,508 - 3180*6.10)* 5

= (18,508 - 19,398)*5

= 4450 UNFAVORABLE